1/x( √4+x^2)^3 using substitution x = 2 tan ϴ

1
Expert's answer
2021-05-19T16:38:09-0400

Substitution


"x=2\tan \theta"

"dx=\dfrac{2d\theta}{\cos^2\theta }"

"\sqrt{4+x^2 }=\sqrt{4+(2\tan \theta)^2 }=\dfrac{2}{\cos\theta }"

"\int\dfrac{dx}{x(\sqrt{4+x^2})^3}=\int\dfrac{2d\theta}{\cos^2\theta (2\tan \theta )(\dfrac{2}{\cos\theta })^3}"

"=\int\dfrac{d\theta}{8\sin \theta }"

"\cos \theta=u"


"du=-\sin \theta d\theta"

"\sin^2 \theta =1-u^2"

"\int\dfrac{d\theta}{8\sin \theta }=-\int\dfrac{ du}{8(1-u^2)}"


"=-\int\dfrac{ du}{16(1-u)}-\int\dfrac{ du}{16(1+u)}"


"=\dfrac{1}{16}\ln|1-u|-\dfrac{1}{16}\ln|1+u|+C"

"=\dfrac{1}{16}\ln(1-\cos\theta)-\dfrac{1}{16}\ln(1+\cos\theta)+C"

"=\dfrac{1}{16}\ln(1-\cos(\arctan\dfrac{x}{2}))"

"-\dfrac{1}{16}\ln(1+\cos(\arctan\dfrac{x}{2}))+C"